OpenStudy (anonymous):

Solve the following system. y = x + 4 5x + y = 16 (2, −6) (−2, 6) (6, 2) (2, 6)

Parth (parthkohli):

You could substitute y with x + 4 in the second equation right there. \( \color{Black}{\Rightarrow 5x + x + 4 = 16}\) Solve for x, and then add 4 to get the value of y.

OpenStudy (shane_b):

All you need to do it plug in the equation for y anywhere you see y in the second equation. Once you do that, you can solve for x in that equation. Using the value of y you found, plug it back into the first equation and solve for y.

OpenStudy (shane_b):

*err I meant "value of x you found" in that last sentence...but you get the idea.

OpenStudy (anonymous):

See I did what yall are saying but I got 1.5 for y and idk how that happened.

OpenStudy (shane_b):

Well, using ParthKohli's equation above, x = 2. You got that far right?

OpenStudy (anonymous):

I did it the way I was though and I took the first equation and made it into y = mx + b to which gave me x = y + 4 then i sub it into the 2nd equation and got 1.5 /:

OpenStudy (shane_b):

You took the wrong off ramp :) This isn't a slope problem...it's a system of equations.

OpenStudy (shane_b):

y=mx+b doesn't have anything to do with this.

OpenStudy (anonymous):

yes b/c the first equation has to be put like that to be substituted, this is the way the teacher taught me

Parth (parthkohli):

Well, you could solve it using the first equation. See where y is 4 more than x, and that'd simply be the answer!

OpenStudy (anonymous):

I got the answer its( 2, 6 ) i plugged it into my graph cal.

Parth (parthkohli):

Yep

Parth (parthkohli):

Well it's the only answer where you see that y is 4 more than x :)

OpenStudy (shane_b):

I would do back-substitution which is what we showed you...that's easier in my mind and is less confusing.

OpenStudy (shane_b):

I don't see the point of trying to put it in slope-intercept form....but that's just me :)

OpenStudy (anonymous):

Yea totally understandable. Idk why she make us do all this when there is easier ways.

OpenStudy (shane_b):

On a test like an SAT you could of course just skip all that and see that y = x +4 basically tells you which answer makes sense

OpenStudy (anonymous):

putting y = x + 4 in 5x + y = 16 6x=16-4 x=2 putting x=2 in y = x + 4 y=6 hope u lyk dis

OpenStudy (anonymous):

(2,6) is right answer. please give me medal if you like my answer!

OpenStudy (anonymous):

To simplify it in simple way, Write first equation is standard form i.e, x + y = c So, you will get, x - y = -4 ---------------1 And the second equation given is: 5x + y = 16 ---------------2 Just add the two equations, x + 5x = -4 + 16 6x = 12 Here solve for x and then put the value of y in any of the above given equation you will have your answer...